OpenStudy (anonymous):

what is 100,000,000 in the power of ten?

OpenStudy (phi):

do you know the power of 10 for 10 ?

OpenStudy (phi):

\( 10^1 \) means 10 multiplied by itself once \( 10^2 \) means 10 multiplied by itself twice \( 10^2 = 100 \)

OpenStudy (phi):

I would look for a pattern \(10^1 = 10 \) power is 1 \(10^2 = 100 \) power is 2 \(10^3 = 1000 \) power is 3 do you see the pattern? the number of zeros happens to equal the power (or exponent)